# Intelligence Test

### v1

@@ 3.3 Classifying Intelligence into Three Levels
In this subsection, we analyze the distribution of failure counts obtained from the Intelligence Test to gain a clear understanding of the subject’s level of intelligence. First, we introduce an Infinity Assumption to define the least intelligent scenario. Based on this, we then propose three levels of intelligence. Finally, we explain how to classify subjects into these three intelligence levels based on the distribution of their failure counts.
@@ 3.3.1 Infinity Assumption
What situation represents a subject having almost no intelligence related to the task? Imagine a scenario in an Intelligence Test where a monkey sits in front of a computer and types to see if it can produce Shakespeare’s works. If it deviates from Shakespeare’s works, we let it attempt again. The failure count refers to the number of attempts before success. The monkey has no understanding of human language and just types randomly. In theory, since the human vocabulary is finite and Shakespeare’s works are also of limited length, the monkey could use an enumeration method, blindly trying all possible combinations of words. Even though most of these combinations are completely nonsensical to us, the monkey can eventually type out Shakespeare’s works. However, this blind, exhaustive enumeration shows that the subject lacks any real intelligence. It is also disconnected from practical reality because the cost of such an exhaustive search would far exceed any reasonable resource limitations, much like how it is completely unrealistic to expect a monkey to eventually produce Shakespeare’s works. Shakespeare did not create his works by randomly typing and waiting for greatness to emerge. Instead, he produced the masterpieces through intentional creativity within the limitations of human life. Therefore, when the failure count approaches the cost of exhaustive enumeration, it almost certainly indicates that the subject has no intelligence related to the task.
We note that the high cost of blind enumeration closely resembles the mathematical concept of infinity. In mathematics, infinity describes a scenario where a quantity is beyond the scale we can measure or endure. For example, when measuring objects on Earth, we can assume the distance from the Sun to the Earth is infinite, and based on this assumption, we treat sunlight as parallel rays. This is because, compared to the size of objects on Earth, the distance between the Sun and Earth is so vast that it can be approximated as infinity. This allows us to use the property of parallel sunlight to help with measurement tasks. The concept of infinity in mathematics is a way of thinking in terms of limits and approximations. While infinity does not directly exist in the physical world, it helps us understand and describe extremely large quantities and allows us to handle them more conveniently. In the case of the Intelligence Test, the characteristic of blind enumeration aligns closely with the concept of infinity. In theory, blind enumeration can eventually lead to the correct solution, but the cost of doing so far exceeds the available resources or our willingness. Therefore, we can model the cost of blind enumeration in the Intelligence Test as infinity and thus can better interpret the results of the test.
@@ 3.3.3 Decay Rate Classification
Before presenting how to practically determine intelligence levels, let us revisit the Infinity Assumption. Although infinity does not exist in the physical world, this does not prevent us from treating certain quantities, which far exceed our capacity to measure or endure, as if they were infinite. In the case of the Intelligence Test, the total number of possible solutions is finite, but as described in the Infinity Assumption, we lack the resources or willingness to blindly enumerate all of them. Therefore, the Infinity Assumption treats the number of possible solutions as if it were infinite.
@@ 3.4 Comparison with Related Work
Since we have introduced Intelligence Test, we can pick up our discussion from Section 2. In contrast to the subjective tests in prior studies, Intelligence Test provides an objective way to evaluate intelligence:
@@
  • Objective Choice of Tasks: We recognize that intelligence is inherently task-dependent. Unlike previous approaches that attempt to define universal intelligence, Intelligence Test does not prescribe any specific task. Instead, it allows researchers to evaluate intelligence in any task of interest, ensuring that the definition of intelligence remains grounded in the actual demands of a given task.
  • Objective Evaluation Framework: The Intelligence Test is mathematically well-defined and does not rely on any hyperparameters. Its conclusions are based on clear statistical criteria rather than subjective assessments. This ensures that evaluations remain consistent across different studies and applications, making it a robust and practical tool for assessing intelligence in real-world settings.
It is important to note that while we argue that intelligence is inherently task-dependent and should be evaluated within specific tasks, this does not prevent researchers from using Intelligence Test as a framework for assessing Artificial General Intelligence (AGI). From our point of view, an AGI system should reach the Autonomous Level in at least every basic human task. Therefore, to evaluate general intelligence, researchers can construct a diverse set of tasks and apply Intelligence Test on each of them.
4 Evaluation with Intelligence Test
In this section, we evaluate state-of-the-art AI systems with Intelligence Test. We adopt a wide range of tasks, including vision, search, recommendation, and language.
@@ 4.1 A Beginner’s Task: MNIST
Figure 2: Experimental Results of Intelligence Test in handwritten digit recognition task (MNIST). The red line is a power law curve drawn based on the distribution of the model’s data points. Results suggest that a system reaches a higher-level of intelligence if it better models the task.
MNIST (Deng, 2012) is a handwritten digit recognition task. It consists of a collection of images depicting the digits 0-9, written by different people. The task is for an AI system to correctly identify the digit in each image. Many people consider MNIST to be a relatively simple task, and it is often used as an introductory challenge for beginners to experiment with various AI algorithms. As such, we also start with this task to test whether Intelligence Test can effectively distinguish different types of AI algorithms.
@@
The experimental results are shown in Figure 2. The blue dots are failure count distribution, and the red line is a power law for reference. The gray, green, and yellow regions represent Limited Level, Capable Level, and Autonomous Level, respectively. As we can see, Intelligence Test clearly distinguishes these three different methods: The linear classifier falls within the Linear Level region, the MLP classifier falls within the Capable Level region, and CNN approaches the Autonomous Level region. Therefore, the better the system models the task, the higher its intelligence level. This suggests that the Intelligence Test is effective at evaluating the intelligence level of an AI system.
@@ 4.2 Vision
Figure 3: Experimental Results of Intelligence Test in Computer Vision. The three rows correspond to three different datasets. Figures in different columns correspond to different models. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results show that models are at Limited Level.
@@
The fact that these models are at Limited Level points to a clear physical meaning: if these models are to guess the answers to a vision-related task, such as finding the correct category or the corresponding images, the model would, in statistical terms, need an infinite number of guesses to get it right. In other words, the model not only makes incorrect predictions but also regards the correct answer as completely wrong. If it tries to solve the task, it will try many incorrect answers before finally outputting the correct one. Therefore, we should not place blind trust in visual models’ results. Instead, we should provide supervision and guidance to ensure their reliability.
@@ 4.3 Search
Next, we evaluate the performance of text search models. Text search should be familiar to many people. It has widespread applications in search engines like Google, Bing, and Baidu. Given a query, the text search model ranks the candidate documents in order of relevance from highest to lowest. We regard this ranking list as its attempt sequence when applying Intelligence Test.
@@
Figure 4: Experimental Results of Intelligence Test in Text Search. The first two rows show performance on the synthetic Wiki task and the web search task. The final row shows performance of the cross-encoder on another four tasks. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results indicate that all models are at the Limited Level.
@@
Limited Level has a clear physical meaning in the text search scenario: when a user submits a query, the document that the user truly wants is likely to be ranked at the end of the list. The user needs to read, in statistical terms, infinite irrelevant documents before reaching the document they are looking for. This highlights the complexity of the text search task and the inadequacy of current search technologies. It inspires us that we cannot simply rely on search engines to seek information.
@@ 4.4 Recommendation
Figure 5: Experimental Results of Intelligence Test in Recommendation System. The first two rows show performance on Product Recommendation and Movie Recommendation. The final row shows the performance of SASRec on another four recommendation tasks. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results indicate that models are at Limited Level.
@@
We adopt many real-world datasets from a wide range of domains. We use the Amazon Beauty dataset (He & McAuley, 2016) to represent e-commerce recommendations. It focuses on skincare product recommendations on the Amazon platform. We use MovieLens (Harper & Konstan, 2015) to represent movie recommendations. It is constructed based on user ratings of movies. We use Steam dataset (Kang & McAuley, 2018) to represent game recommendations. It recommends games to players on the Steam platform. We use Douban Book (Zhu et al., 2020, 2019) to represent book recommendations. Douban is a popular Chinese internet platform and this dataset is to recommend books to users. We use Douban Music (Zhu et al., 2020, 2019) to represent music recommendations. It is also collected from the Douban platform and is to recommend music to users. Finally, we use Gowalla dataset (Cho et al., 2011) to represent location recommendations. Gowalla is a location-based online social network application where users share their current location. The dataset is to recommend places users might like to visit.
We test four widely recognized recommendation methods. The first is a popularity-based recommendation method. As the name suggests, it ranks items based on their popularity and recommends them accordingly. Although it is straightforward, it is effective and commonly used in real-world applications. The other three methods are neural-based models: GRU4Rec (Hidasi et al., 2016; Hidasi & Karatzoglou, 2018), SASRec (Kang & McAuley, 2018), and ComiRec (Cen et al., 2020). They differ in architecture. GRU4Rec employs recurrent neural networks to build user profiles based on the interaction history. SASRec uses attention mechanisms to model how past interactions influence future preferences. ComiRec captures users’ diverse interests with a dynamic modeling approach.
@@
This result has a clear physical meaning: in current recommendation applications, users will see, in statistical terms, infinite uninterested items before being presented with something they are truly interested in. If a user is disappointed every time they see an item they are not interested in, it means the current recommendation system will disappoint them countless times.
@@ 4.5 Language
We have assessed AI models in vision, search, and recommendation tasks. Now, we proceed to language tasks. Some studies claim that large language models have already achieved exceptionally high-level intelligence and passed the Turing Test (Biever, 2023; Aharoni et al., 2024; Mei et al., 2024). With Intelligence Test, we can examine their intelligence levels and re-think this conclusion. We will use four tasks for a comprehensive evaluation, including coding, mathematics, question answering, and writing.
@@ 4.5.1 Coding
Figure 6: Experimental Results of Intelligence Test in Coding. The three rows correspond to three datasets, and figures in different columns correspond to different models. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. The results indicate that models are approaching Capable Level.
@@
The experimental results are shown in Figure 6. The three rows present results on HumanEval, MBPP, and CRUXEval, respectively. We can see that models with more parameters are closer to the Capable Level. For 70B models, a few data points are already within the Capable Level region, yet a long tail of data points still falls at the Limited Level region. Therefore, although current models are relatively strong and approaching the Capable Level in coding, they are mostly at the Limited Level. It means that they cannot reliably find correct solutions for basic coding questions. Thus, human supervision is essential.

### v2

@@ 3.3 Classifying Intelligence into Three Levels
In this subsection, we analyze the distribution of failure counts obtained from the Survival Game to gain a clear understanding of the subject’s level of intelligence. First, we introduce an Infinity Assumption to define the least intelligent scenario. Based on this, we then propose three levels of intelligence. Finally, we explain how to classify subjects into these three intelligence levels based on the distribution of their failure counts.
@@ 3.3.1 Infinity Assumption
What situation represents a subject having almost no intelligence related to the task? Imagine a scenario in an Survival Game where a monkey sits in front of a computer and types to see if it can produce Shakespeare’s works. If it deviates from Shakespeare’s works, we let it attempt again. The failure count refers to the number of attempts before success. The monkey has no understanding of human language and just types randomly. In theory, since the human vocabulary is finite and Shakespeare’s works are also of limited length, the monkey could use an enumeration method, blindly trying all possible combinations of words. Even though most of these combinations are completely nonsensical to us, the monkey can eventually type out Shakespeare’s works. However, this blind, exhaustive enumeration shows that the subject lacks any real intelligence. It is also disconnected from practical reality because the cost of such an exhaustive search would far exceed any reasonable resource limitations, much like how it is completely unrealistic to expect a monkey to eventually produce Shakespeare’s works. Shakespeare did not create his works by randomly typing and waiting for greatness to emerge. Instead, he produced the masterpieces through intentional creativity within the limitations of human life. Therefore, when the failure count approaches the cost of exhaustive enumeration, it almost certainly indicates that the subject has no intelligence related to the task.
We note that the high cost of blind enumeration closely resembles the mathematical concept of infinity. In mathematics, infinity describes a scenario where a quantity is beyond the scale we can measure or endure. For example, when measuring objects on Earth, we can assume the distance from the Sun to the Earth is infinite, and based on this assumption, we treat sunlight as parallel rays. This is because, compared to the size of objects on Earth, the distance between the Sun and Earth is so vast that it can be approximated as infinity. This allows us to use the property of parallel sunlight to help with measurement tasks. The concept of infinity in mathematics is a way of thinking in terms of limits and approximations. While infinity does not directly exist in the physical world, it helps us understand and describe extremely large quantities and allows us to handle them more conveniently. In the case of the Survival Game, the characteristic of blind enumeration aligns closely with the concept of infinity. In theory, blind enumeration can eventually lead to the correct solution, but the cost of doing so far exceeds the available resources or our willingness. Therefore, we can model the cost of blind enumeration in the Survival Game as infinity and thus can better interpret the results of the test.
@@ 3.3.3 Decay Rate Classification
Before presenting how to practically determine intelligence levels, let us revisit the Infinity Assumption. Although infinity does not exist in the physical world, this does not prevent us from treating certain quantities, which far exceed our capacity to measure or endure, as if they were infinite. In the case of the Survival Game, the total number of possible solutions is finite, but as described in the Infinity Assumption, we lack the resources or willingness to blindly enumerate all of them. Therefore, the Infinity Assumption treats the number of possible solutions as if it were infinite.
@@
3.4 Approximation with Reference Answers
Note that Survival Game requires to determine whether each attempt made by a subject is correct. Yet verifying correctness for every attempt can be expensive in some tasks. Consider a task where the test subject is to prove a mathematical theorem. The subject provides proof with each attempt. However, for complex mathematical theorems, the proofs can be very long and intricate, and the cost of verifying the correctness of each proof is extremely high. In such cases, directly applying Survival Game may make the process prohibitively expensive.
Therefore, we propose an approximation method to address this problem. We will first introduce the underlying assumption and then formalize the approximation method.
3.4.1 Scale-Invariance Assumption
In situations where it is difficult to verify the correctness of each trial, we can adopt an alternative approach: counting the failure attempts before the subject arrives at a predefined reference answer. For example, when testing whether a subject can prove mathematical theorems, we do not evaluate whether each of its outputs constitutes a valid new proof. Instead, we check whether it can produce a known proof.
To support the validity of such reference-based evaluation, we propose a Scale-Invariance Assumption:
  • The number of failures before finding any solution follows a power-law distribution.
  • The number of failures before finding any solution is linearly related to the number of failures before finding a particular solution.
Under the given assumption, we can prove that the number of failures before finding any solution and the number of failures before finding a particular solution have the same failure decay rate. Therefore, they yield the same result in our Decay Rate Classification. The proof is as follows:
$P(X=x)=Cx^{-\alpha},\quad x\geq x_{\text{min}}.$ (1)
$P(kX=x)=P(X=x/k)=C(x/k)^{-\alpha}=Ck^{\alpha}x^{-\alpha}\propto x^{-\alpha}.$ (2)
Thus, the power-law distribution retains its functional form under linear scaling. If the failure count follows the power law, a linear transformation does not change the power-law formulation and the exponent.
The first part of the Scale-Invariance Assumption is empirically and theoretically supported. Specifically, Section 5 will show that the failure count is close to a power-law distribution. In Section 6, we theoretically analyze the cause of such a phenomenon. We demonstrate that this is because human tasks exhibit criticality property.
The second part of the Scale-Invariance Assumption requires further investigation. Whether it is linearly correlated depends on the relationship between references, tasks, and subjects. In our experiments, we use human-written answers as references. For example, when assessing whether the model can write mathematical proofs, we use human-written proofs as references. When evaluating whether the model can generate high-quality legal opinions, we use legal opinions written by human judges. Similarly, when assessing whether the model can produce excellent literary works, we use human literary works as references. In these cases, we assume that the number of failed attempts to arrive at a feasible solution is linearly related to the number of attempts to reach these reference solutions. We have not verified its correctness for now and will verify it in the future.
3.4.2 Survival Game with References
Based on the Scale-Invariance Assumption, we propose a variation named Survival Game with References. It avoids the need for direct correctness verification while keeping the core of Survival Game. In those tasks where the cost of correctness verification is high, it uses a reference answer and measures the number of failed attempts before producing the reference answer. The validity of this method is supported by the following theorem:
Theorem 3.4 (Survival Game with References)
Let $X^{*}$ be a discrete random variable representing the number of failure attempts before finding a predefined reference answer. $X^{*}$ is an upper bound estimation of the real failure counts. If the Scale-Invariance Assumption holds, the failure decay rate of $X^{*}$ is accurate.
This approach eliminates the need for verifying every attempt and instead examines failure counts until reaching a known reference answer. It is a low-cost realization of Survival Game and reflects an upper bound of the subjects’ errors. If the Scale-Invariance Assumption holds, this theorem shows that we can exactly evaluate the intelligence level in an efficient way.
@@ 3.5 Comparison with Related Work
Since we have introduced Survival Game, we can pick up our discussion from Section 2. In contrast to the subjective tests in prior studies, Survival Game provides an objective way to evaluate intelligence:
@@
  • Objective Choice of Tasks: We recognize that intelligence is inherently task-dependent. Unlike previous approaches that attempt to define universal intelligence, Survival Game does not prescribe any specific task. Instead, it allows researchers to evaluate intelligence in any task of interest, ensuring that the definition of intelligence remains grounded in the actual demands of a given task.
  • Objective Evaluation Framework: The Survival Game is mathematically well-defined and does not rely on any hyperparameters. Its conclusions are based on clear statistical criteria rather than subjective assessments. This ensures that evaluations remain consistent across different studies and applications, making it a robust and practical tool for assessing intelligence in real-world settings.
It is important to note that while we argue that intelligence is inherently task-dependent and should be evaluated within specific tasks, this does not prevent researchers from using Survival Game as a framework for assessing Artificial General Intelligence (AGI). From our point of view, an AGI system should reach the Autonomous Level in at least every basic human task. Therefore, to evaluate general intelligence, researchers can construct a diverse set of tasks and apply Survival Game on each of them.
4 Evaluation with Survival Game
In this section, we evaluate state-of-the-art AI systems with Survival Game. We adopt a wide range of tasks, including vision, search, recommendation, and language.
@@ 4.1 A Beginner’s Task: MNIST
Figure 3: Experimental Results of Survival Game in handwritten digit recognition task (MNIST). The red line is a power law curve drawn based on the distribution of the model’s data points. Results suggest that a system reaches a higher-level of intelligence if it better models the task.
MNIST (Deng, 2012) is a handwritten digit recognition task. It consists of a collection of images depicting the digits 0-9, written by different people. The task is for an AI system to correctly identify the digit in each image. Many people consider MNIST to be a relatively simple task, and it is often used as an introductory challenge for beginners to experiment with various AI algorithms. As such, we also start with this task to test whether Survival Game can effectively distinguish different types of AI algorithms.
@@
The experimental results are shown in Figure 3. The blue dots are failure count distribution, and the red line is a power law for reference. The gray, green, and yellow regions represent Limited Level, Capable Level, and Autonomous Level, respectively. As we can see, Survival Game clearly distinguishes these three different methods: The linear classifier falls within the Linear Level region, the MLP classifier falls within the Capable Level region, and CNN approaches the Autonomous Level region. Therefore, the better the system models the task, the higher its intelligence level. This suggests that the Survival Game is effective at evaluating the intelligence level of an AI system.
@@ 4.2 Vision
Figure 4: Experimental Results of Survival Game in Computer Vision. The three rows correspond to three different datasets. Figures in different columns correspond to different models. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results show that models are at Limited Level.
@@
The fact that these models are at Limited Level points to a clear physical meaning: if these models are to find out the answers to a vision-related task when they are wrong in the first place, the model would, in statistical terms, need an infinite number of guesses to get it right. In other words, the model not only makes incorrect predictions but also regards the correct answer as completely wrong. If it tries to solve the task, it will try many incorrect answers before finally outputting the correct one. Therefore, we should not place blind trust in visual models’ results. Instead, we should provide supervision and guidance to ensure their reliability.
@@ 4.3 Search
Next, we evaluate the performance of text search models. Text search should be familiar to many people. It has widespread applications in search engines like Google, Bing, and Baidu. Given a query, the text search model ranks the candidate documents in order of relevance from highest to lowest. We regard this ranking list as its attempt sequence when applying Survival Game.
@@
Figure 5: Experimental Results of Survival Game in Text Search. The first two rows show performance on the synthetic Wiki task and the web search task. The final row shows performance of the cross-encoder on another four tasks. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results indicate that all models are at the Limited Level.
@@
Limited Level has a clear physical meaning in the text search scenario: when a user submits a query and the right document is not ranked at the top, the right document is likely to be ranked at the end of the list. The user needs to read, in statistical terms, infinite irrelevant documents before reaching the document they are looking for. In other words, when a search model makes a mistake, it is almost completely unable to correct itself. This highlights the complexity of the text search task and the inadequacy of current search technologies. It inspires us that we cannot simply rely on search engines to seek information.
@@ 4.4 Recommendation
Figure 6: Experimental Results of Survival Game in Recommendation System. The first two rows show performance on Product Recommendation and Movie Recommendation. The final row shows the performance of SASRec on another four recommendation tasks. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results indicate that models are at Limited Level.
@@
We adopt many real-world datasets from a wide range of domains. We use the Amazon Beauty dataset (He & McAuley, 2016) to represent users’ preference in e-commerce recommendations. It focuses on skincare product recommendations on the Amazon platform. We use MovieLens (Harper & Konstan, 2015) to represent movie recommendations. It is constructed based on user ratings of movies. We use Steam dataset (Kang & McAuley, 2018) to represent users’ preference in game recommendations. It recommends games to players on the Steam platform. We use Douban Book (Zhu et al., 2020, 2019) to represent book recommendations. Douban is a popular Chinese internet platform and this dataset is to recommend books to users. We use Douban Music (Zhu et al., 2020, 2019) to represent users’ preference in music recommendations. It is also collected from the Douban platform and is to recommend music to users. Finally, we use Gowalla dataset (Cho et al., 2011) to represent location recommendations. Gowalla is a location-based online social network application where users share their check-in location. The dataset is to recommend places users might like to visit.
We test four widely recognized recommendation methods. The first is a popularity-based recommendation method. As the name suggests, it ranks items based on their popularity and recommends them accordingly. Although it is straightforward, it is effective and commonly used in real-world applications. The other three methods are sequential recommendation models: GRU4Rec (Hidasi et al., 2016; Hidasi & Karatzoglou, 2018), SASRec (Kang & McAuley, 2018), and ComiRec (Cen et al., 2020). They differ in architecture. GRU4Rec employs recurrent neural networks to build user profiles based on the interaction history. SASRec uses attention mechanisms to model how past interactions influence future preferences. ComiRec captures users’ diverse interests with a dynamic modeling approach.
@@
This result has a clear physical meaning: When a recommendation model makes a mistake, it is almost impossible for it to find the correct product through continuous attempts. For users, it means that users will see, in statistical terms, infinite uninterested items before being presented with something they are truly interested in. If a user is disappointed every time they see an item they are not interested in, the current recommendation system will disappoint them countless times.
@@ 4.5 Language
We have assessed AI models in vision, search, and recommendation tasks. Now, we proceed to language tasks. Some studies claim that large language models have already achieved exceptionally high-level intelligence and passed the Turing Test (Biever, 2023; Aharoni et al., 2024; Mei et al., 2024). With Survival Game, we can examine their intelligence levels and re-think this conclusion. We will use four tasks for a comprehensive evaluation, including coding, mathematics, question answering, and writing.
@@ 4.5.1 Coding
Figure 7: Experimental Results of Survival Game in Coding. The three rows correspond to three datasets, and figures in different columns correspond to different models. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. The results indicate that models are approaching Capable Level.
@@
The experimental results are shown in Figure 7. The three rows present results on HumanEval, MBPP, and CRUXEval, respectively. We can see that models with more parameters are closer to the Capable Level. For 70B models, a few data points are already within the Capable Level region, yet a long tail of data points still falls at the Limited Level region. Therefore, although current models are relatively strong and approaching the Capable Level in coding, they are mostly at the Limited Level. It means that they cannot reliably find correct solutions through trial and error for basic coding questions. Thus, human supervision is essential.